Maria mixed 4% and 8% acid solutions to obtain a mixture of 500 mL which is 6.4% acid. How much
of the 4% and 8% solution were used?

Answer by addingup(3677)

mixed 4% and 8% acid solutions to obtain a mixture of 500 mL which is 6.4%
Let the 4% solution be x
0.04x+0.08(500-x) = 0.064(500)
0.04x+40-0.08x = 32 subtract 40 on both sides and subtract on left
0.04x-0.08x+40-40 = 32-40
-0.04x = -8 divide both sides by -0.04 and remember -/- = +
x = 200
Maria used 200mL of the 4% solution and 300mL of the 8% solution